Sol Summary Report – February 9th

Crew 188 Sol 12 Summary Report 9FEB2018

Sol 12

Summary Title: One more round then it’s [oxygen] bottles to the ground

Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Mission Status: EVAs are complete, it’s the final crunch.

Sol Activity Summary:

As the Olympic flame is lit back on Earth, our physical challenges in the field come to a close. With two final EVAs, the crew is dusting off the suits and preparing them for the next crew who will start a week later. The morning EVA hiked up to the Hab Ridge for some final 360 filming, chlorophyll sampling and scenic views. The afternoon EVA had a few samples along the way to Toothy Ridge to the North. Final working photos were snapped with a few last planned for tomorrow during our campus cleanup.

As with any online event on Earth, our Mars-2-Mars event will have a slight change of plans with the session being recorded on WebEx and shared immediately afterwards by the Austrian Space Forum (Österreichisches Weltraum Forum, OeWF). Q&A can be submitted to reddit and the crews will address them during the hangout or afterwards as an ongoing discussion.

10Feb2018, 17:00 CET = 09:00 MDRS time (Sol 13) = 20:00 Oman time [16:00 GMT]

reddit Q&A: https://www.reddit.com/r/Mars/comments/7s23yz/mdrs_crew_188_ama

Spacesuit Up!

Ryan L. Kobrick, Ph.D.

MDRS Crew 188 Commander

Look Ahead Plan:

A hangout with #SimulateMars in Oman, the AMADEE-18 mission at Kepler Station. Then we will end the MDRS Crew 188 simulation mid-day as we continue to clean all areas. With our continuous Hab-chores cleaning schedule we should be on track for an early Sunday morning departure.

Anomalies in work:

None.

Weather:

Calm winds and hazy skies with warm dropping all day.

Crew Physical Status:

There is a mixed of anxiety to get work completed, excitement to return home to loved ones, and sadness to be separating form our Martian family.

EVA:

15 EVAs were completed during our mission.

Reports to file:

1. Ops Report

2. Sol Summary

3. Journalist Report

4. Green Hab Report

5. EVA #14 and #15 report

6. Daily Photos

7. Health & Safety Officer Report

8. Mission Summary

Support Requested (same as previous three reports):

– Heater for science dome requested. A purchase order request was made.

– FYI for mission support to be aware of for systems to track and safe operations.

– There are potentially two packages at Hollow Mountain that we would love to have picked up.

– We will pick them up post-sim

– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open.

Sol Summary – February 7th

Crew 188 Sol 10 Summary Report 7FEB2018

Sol 10

Summary Title: Game On!
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: Crew feeling good after a full day.

Sol Activity Summary:
Even with one EVA, the crew had a very busy and productive day. The EVA challenged the crew to a new slew of Martian Olympiad events counting down to the Winter Olympic Games back on Earth. An opening ceremony occurred and the crew hopes to share this artistic footage in high resolution 360 views. There are sure to be phots posted across social media, just look for #MDRS188. Meanwhile in the Hab, the NASA Glenn dust load experiment continued and air filtration dust samples were obtained from the EVA crew returning into the airlock. The afternoon was highlighted with staring at the Sun with a Hydrogen-alpha filter seeing sunspots, flares and even an airplane captured on camera passing in front of the Sun! The crew was busy digitally capturing video footage and audio for podcasts, school groups and for a mission documentary. After sunset the crew enjoyed watching the International Space Station pass-by briefly and an Iridium satellite. Who knows what space infrastructure Mars inhabitants will be able to watch from their homes, but they will need some form of planetary connectivity.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
Two EVAs are panned that will fill most of the operational day. Afterwards the third and final human factors group chat will occur.

Anomalies in work:
Propane was detected in the afternoon. The crew is carefully monitoring and performing basic bubble testing on pipes. Currently no leaks are detected.

Weather:
Calm winds and clear skies with temperatures dropping rapidly at sunset.

Crew Physical Status:
One crewmember has cold symptoms, which may be due to fatigue but health status is nominal.

EVA:
Two are planned for tomorrow.

Reports to file:
1. Ops Report
2. Sol Summary
3. Journalist Report (combined Sol 9 and 10)
4. Green Hab Report
5. EVA #13 and #14 request
6. EVA #12 Reports
7. Daily Photos

Support Requested (same as previous report):
– Heater for science dome requested. A purchase order request was made.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open. [end]

Sol Summary – February 6th

Crew 188 Sol 9 Summary Report 6FEB2018

Sol 9

Summary Title: Sacrifice
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: Crew is well.

Sol Activity Summary:

I think it goes without saying that our crew made many sacrifices to be a part of this two-week experience (finances, family, work, and much more). One that may not be obvious, and would have been impossible to time, was missing the Falcon Heavy historic launch. The unquestionable highlight for today was SpaceX’s success launch and landings, and what that may mean for future human spaceflight exploration. Today was packed with exploration. Two EVAs including a beautifully choreographed performance optimizing a 360 camera, and a driving excursion to uncharted territory. The crew was able to watch the launch and twin landings in a video later in the evening. Wow. Even watching the launch delayed had a similar impact to being there with family and friends in Florida. It was rocket-eye-candy-overload. Who knows when the next FH launch will be, or where any of us may be located, but we do know that these are the ingredients needed to make a mission to Mars feasible.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
Tomorrow solar observations will continue in the morning hoping for calmer weather. An afternoon EVA will celebrate our international diversity and the upcoming Olympic Games with our own Martian Olympiad.

Anomalies in work:
None.

Weather:
The temperatures are cooling off and the winds remain high.

Crew Physical Status:
The crew has been burning on full throttle from 8am until 11pm and pushing hard to the finish line to be fully successful in their scientific and artistic investigations at MDRS.

EVA:
One is planned for tomorrow.

Reports to file:
1. Ops Report
2. Sol Summary
3. Journalist Report
4. Green Hab Report
5. EVA #12 request
6. EVA #10 and #11 Reports
7. Daily Photos

Support Requested:
– Heater for science dome requested. A purchase order request was made.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open.
– Our reports have not been posted to the MDRS website since Saturday Feb 3rd. Please check the status of these documents for dissemination. We did see some photos on Facebook.

Sol Summary – February 5th

Crew 188 Sol 8 Summary Report 05FEB2018

Sol 8

Summary Title: Team Unity
Author’s name: Renee Garifi, MDRS Crew 188 Executive Officer Mission Status: Nominal

Sol Activity Summary:
This morning our team awoke to a beautifully and thoughtfully arranged breakfast table with a special greeting written by two crewmembers who wanted to help our team start week 2 of our mission on a high note. Each mission brings with it a unique set of challenges and rewards and I am humbled and proud to be part of such a courageous and strong team.

Every morning at 08:30 we hold our Daily Planning Conference and one crewmember hosts a briefing of the upcoming EVA or team activity that will require the most coordination by all crew. We spent the mid-morning preparing for a 4-person EVA with the objective of scouting filming locations for a performing astronautics project by our Artist in Residence utilizing the Insta360 Pro camera. Two of our crew remained behind at the Hab to serve as HABCOM on the radio with the EVA team and keep up with science projects and tasks around the Hab.

Following the EVA we shared lunch and set about working on our individual projects. We had a productive afternoon followed by a period of relaxation. Our crew is currently in good spirits and communication, stress management and team work is at an all-time high. In the evenings when our crew is not too exhausted, we have started hosting a lecture series here on Mars. After our comms window closes and our crew finishes dinner and dishes, the next scheduled crewmember to present will spend about 30-45 minutes sharing a topic that they are either passionate about or teaching a useful skill. Our crew is comprised of six individuals with strong personal and professional backgrounds that make for informative evening sessions that bond and inspire us.

Ad Astra!
Renee Garifi
MDRS Crew 188 Executive Officer

Look Ahead Plan:
We will perform EVA #10 in the morning tomorrow and EVA #11 in the afternoon.

Anomalies in work:
None.

Weather:
Windy, but not like the high winds of yesterday. Skies were cloudy.

Crew Physical Status:
The crew is tired but in good spirits.

EVA:
Two EVAs planned for tomorrow.

Reports to file:
1. EVAs #10 and #11 request
2. EVA #9 Report
3. Ops Report
4. Sol Summary
5. Journalist Report
6. Green Hab Report
7. Daily Photos

Support Requested (acknowledged yesterday but listed as reminders): – Heater for science dome requested.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Water tank refills will be needed in the next few days.

Sol Summary Report – February 4th

Crew 188 Sol 7 Summary Report 4FEB2018

Sol 7

Summary Title: Turning Point

Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander

Mission Status: Crew is well.

Sol Activity Summary:

Today’s highlights included a pedestrian EVA back to Hab Ridge for micrometeorite and chlorophyll sample collection and scenic views with Tat, Julia, and Renee. The crew enjoyed performing the bubbles experiment a few times including by the science dome for a short video recording. During the morning, Sarah Jane and Zac helped with the spacesuit mobility study that I am conducting with my ERAU S.U.I.T. Lab (see @SpacesuitUp) at several analogue station locations around the world. SJ worked as the technician and Zac the suited/unsuited test subject.

Afterwards the crew had a big lunch followed by round 2 of Tat’s human factors study. It was a very enlightening conversation that identified how we could all help each other with our scientific goals during the mission. There are a lot of projects underway being managed in a delicate balance and a short time budget. Many of the projects already have publication plans lined up at peer reviewed conferences or with partnerships at universities and museums for exhibits. We owe it to our partner PIs around the world to extract the best possible data possible while roaming the red desert.

In the early evening, our #SuperScienceSunday blitz did not go as planned to help reach the public with our messages about exploration and education. Not all plans happen as intended, but flexibility and perseverance will keep our team bonded to face challenges to share something bigger, something that cannot be explained in abundant daily reports and occasional snaps on social media. Our crew wound down with a group meditation session in the science dome, followed by open and very personal discussion, and was crowned with a yoga session to stretch our legs for the next EVA and the next challenge.

Spacesuit Up!

Ryan L. Kobrick, Ph.D.

MDRS Crew 188 Commander

Look Ahead Plan:

Tomorrow EVA 9 will take the crew to a site first explored by Crew 187 in the south regions of our exploration zone. Crew members will work on various projects at the Hab. The following day is likely to be followed by two EVAs so that we can meet our exploration and scientific objectives. Make sure to check out updates on Facebook: https://www.facebook.com/TeamISUonMars and Twitter: https://twitter.com/ISUonMars.

Anomalies in work:

None.

Weather:

Windy, but not like the high winds of yesterday. Skies were mostly clear.

Crew Physical Status:

The crew is well, but is enjoying a bit of personal space as well.

EVA:

One is planned for tomorrow.

Reports to file:

1. EVA #9 request

2. EVA #8 Report

3. Ops Report

4. Sol Summary

5. Journalist Report

6. Green Hab Report

7. Daily Photos

Support Requested (acknowledged yesterday but listed as reminders):

Heater for science dome requested.

There are potentially two packages at Hollow Mountain that we would love to have picked up.

Water tank refills will be needed in the next few days.

Sol Summary – February 3rd

Sol 6

Summary Title: A Blustery Day on Mars
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander
Mission Status: Everything is fine, how are you?

Sol Activity Summary:
A two-EVA day today kept the crew busy all day. The morning EVA (Renee, Julia, and I) went North to discover the "Goblins" in Lith Canyon. The crew followed an ancient stream bed into the canyon, a familiar site for myself once we reached a few of the more memorable dips, with a grand opening from rolling red cliffs to an open field followed by a deep canyon drop area near what we hope to confirm was the correct location.

The afternoon EVA (Zac, Tat, and SJ) can be summarized by two words, windy bubbles. EVA double-oh-seven ventured by foot North and scrambled up the red cliffs near Sagan Road to reach the Hab Ridge Road. Performing a variety of bubbles experiments, the crew experience upper wind limits for their expedition. They walked south towards the Hab for a scenic view of our campus from above.

Meanwhile back in the Hab, I was in full science mode taking dust measurements in the airlock and uploading biometric data to Earth. The winds are expected to keep the Hab cool overnight leading into Super Science Sunday… I mean Super Bowl Sunday. The crew is planning to coordinate outreach activities to see if we can help raise awareness for human spaceflight and Mars-forward research.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
EVA 8 is being planned to take the crew back up on the Hab Ridge for sample collection and scenic views. The ERAU spacesuit mobility study is being planned for video capture in the science dome. IN the afternoon we will have our next installment of the human factors study. During the day the team will be trying to post awesome social media content for #SuperScienceSunday.

Have a better hashtag?

Let us know on Facebook: https://www.facebook.com/TeamISUonMars and Twitter: https://twitter.com/ISUonMars.

Anomalies in work:
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
A very windy day. Skies were mostly clear.

Crew Physical Status:
The crew is in the groove zone.

EVA:
One is planned for tomorrow.

Reports to file:
1. EVA #8 request
2. EVA #6 and #7 Report
3. Ops Report
4. Sol Summary
5. Journalist Report
6. Green Hab Report
7. Daily Photos

Support Requested:
Heater for science dome requested.
There are potentially two packages at Hollow Mountain that we would love to have picked up.
Water tank refills will be needed in the next few days.

Sol Summary Report – February 2nd

Crew 188 Sol 5 Summary Report 2FEB2018

Sol 5

Summary Title: Mars Rocks!

Author’s name: Renee Garifi, MDRS Crew 188 Executive Officer

Status: Almost half-way!

Sol Activity Summary:

Today the crew completed EVA #5 and had a successful afternoon of setting up scientific research hardware, working on personal outreach projects and getting to the overdue tasks that have been previously placed on hold in order to keep pace with the busy mission. We will soon wrap up week number 1 in our home sweet Hab here at MDRS.

The EVA team was excited to share their geological samples collected in the field with the HAB team as soon as they arrived back through the airlock. Julia’s favorite rock found today was a vibrant eye-catching red color that seemed otherworldly and intriguing. They explored a canyon route along Cactus Road and took breathtaking photos and videos of their trek while scouting sights for future science sample collections. Tomorrow, we will be conducting two EVAs with the first scouting for potential micrometeorite deposit sights and collect top soil samples from hill tops to be analyzed for very small micrometeorite rocks. The second will have the crew exploring Hab Ridge by foot to continue to explore our surroundings.

The crew took some well-deserved personal time this afternoon to refresh and gear up for closing out week 1 of our mission. We are looking forward to watching Groundhog Day for the humor and irony (today is Groundhog Day in the USA).

Ad Astra!

Renee Garifi

MDRS Crew 188 Executive Officer

Look Ahead Plan:

EVA 6 is planned for tomorrow to revisit a site from EVA 2 and to further add navigation and waypoint data for the crew to use on all EVAs. EVA 7 is a pedestrian EVA to Hab Ridge. Meanwhile several other projects will be underway inside the dome.

Anomalies in work:

Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:

The weather was clear today with few clouds.

Crew Physical Status:

It was a busy day all around, so we are looking at our mission ahead to plan out key EVAs, science projects, and fun activities to make sure we are having a well-balanced experience.

EVA:

One today, Two tomorrow.

Reports to file:

1. EVA #5 Report

2. EVA #6 and #7 Request

2. Ops Report

3. Sol Summary

4. Journalist Report

5. Creative Report

6. Green Hab Report

7. Daily Photos

Support Requested: Our crew’s reports have not been posted to the MDRS website yet. We would like to request that these be back-posted so that all of our reports and the work we have put into the mission get reflected on the MDRS website.

Per direction from Dr. Rupert, we have removed the broken microwave from the Hab and placed it into the back of the DEIMOS rover for removal. We anxiously await the arrival of our replacement unit by cargo delivery ship.

Sol Summary – February 1st

Crew 188 Sol 4 Summary Report 1FEB2018

Sol 4

Summary Title: A sol featuring soul with the devastation of soles Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: All nominal

Sol Activity Summary:
Let’s jump to the bottom line. My hiking boots were completely destroyed on our hike to Phobos Peak with the soles peeling off the bottom. An EVA with modest to difficult objectives was challenged with equipment failures and ingenuity to solve problems as they occurred. This is the type of exploration we need for a Mars mission. Calculating risks, making decisions, and with collective exploration experience solving problems. There’s going to be a lot of problems. The EVA Macgyver of the day was Sarah Jane, fixing one boot with a safety whistle, and the other with my laces and some zip ties I was using. Between Julia, SJ and I we were challenged physically and mentally. I had suggested Phobos Peak for the EVA, and I was clearly willing to lose my boots over it. Luckily my hiking shoes are rugged and ready to step up to the next EVA challenge.

The crew enjoyed the scenic views from the science dome during a session of Kung Fu Thai Chi in the morning. Every morning needs to start off on the right foot… even if part of your foot falls off later. The crew also cracked into Cottage Cheese with Blackcurrant Puree space desert from the Russian Space Food Laboratory. It’s apparently “approved by space”. Further work continued on the NASA VEGGIE OASYS project in the Green Hab. We also finally opened the dust load experiment that I am running from NASA Glenn Research Center. I’m excited to set up the hardware and see how much dust we are tracking into the hab. It’s a dusty environment, and knowing how dirty it gets during key activities (EVA airlock, post EVA clean up, and more) we can help prepare future life support systems with the correct type of filters.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
EVA 5 is planned for tomorrow to revisit a site from EVA 2 and to further add navigation and waypoint data for the crew to use on all EVAs. Meanwhile several other projects will be underway inside the dome.

Anomalies in work:
MDRS Director is coordinating power system updates with contractors. There was a site visit today.
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
The weather was clear until mid-day when clouds rolled into the area.

Crew Physical Status:
It was a busy day all around, so we are looking at our mission ahead to plan out key EVAs, science projects, and fun activities to make sure we are having a well-balanced experience.

EVA:
One is on the books for tomorrow.

Reports to file:
1. EVA #5 requests
2. Ops Report
3. Sol Summary
4. Journalist Report
5. Creative Report
6. Green Hab Report
7. Daily Photos

Support Requested:
Website updates.
Microwave timer works, but main chamber is not functioning/heating. [end]

Sol Summary – January 31st

Crew 188 Sol 3 Summary Report 31JAN2018

Sol 3

Summary Title: Winding down before ramping up
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: All nominal

Sol Activity Summary:
Sol 2 merged into Sol 3 with the astronomical Super Blue Blood Moon event! The crew had varying levels of sleep hours leading into the early morning event. As you may have seen from earlier reports this was a very rare event, and the show met all expectations. All crew members were able to watch the Moon slowly disappear behind Earth’s shadow leaving a red glowing Moon. Looking at the red Moon created nice mental links to our Martian experience and submersion and let us ponder what it would be like to watch a dual-lunar event on Mars. Because of the rare event, today was dedicated to ramping up our science projects and resting for future EVAs. Personally I worked on my EVA metric study with the help of Tat to edit and name waypoints and record key data into a master spreadsheet. I also recorded a video podcast to post publicly (hopefully soon) interviewing Dr. Sarah Jane Pell. I am sure my students will find it very interesting and I’m expecting their follow up questions will be great! Renee and Zac started working on setting up the prototype of NASA Kennedy Space Center’s VEGGIE OASYS experiment in the Green Hab. It is great that the crew gets to work on an experiment that may one day be used on Mars, and super cool that it is already on the International Space Station. Thank you Dr. Luke Roberson for being the quarterback of this project at NASA KSC! Julia and Zac were also working in the Green Hab today, optimizing the layout and tending to the plants. Tat started a human factors analysis for the MDRS operational structure. This study could lead to consolidated task management and more efficient processes. The lighter day allowed the crew to sync files and prepare to share their experience. Several crewmembers sewed their mission patches onto their flight suits today, preparing to explore in style.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
In the morning three crew members will head East to Phobos peak to examine human performance and capture 360 footage in 8k. The afternoon EVA will scout more roads and re-visit a previous site. The crew will be fairly busy with these two EVAs, but other projects will continue to unfold.

Anomalies in work:
MDRS Director is coordinating power system updates with contractors. There was a site visit today.
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
The weather was clear all day and provided great views out the surrounding terrain.

Crew Physical Status:
Hydration and altitude adjustment seems to be going well. Some minor skin dryness for some crewmembers. Everyone seems more rested after a more relaxed day on campus.

EVA:
Two EVAs are being planned and requested for tomorrow.

Reports to file:
1. EVA #4 and #5 requests
2. Ops Report
3. Sol Summary
4. Journalist Report
5. Creative Report
6. Green Hab Report
7. Astronomy Report
8. Daily Photos

Support Requested:
None.

Sol Summary – January 30th

Crew 188 Sol Summary Report 30JAN2018

Sol 2

Summary Title: The Human Factor
Author’s name: Renee Garifi, MDRS Crew 188 Executive Officer Mission Status: On track

Sol Activity Summary:

We awoke this morning to a beautiful Martian sunrise and amazing view of the alien landscape out the port hole windows on the second floor of the hab. We held our daily planning conference over breakfast where we mapped out our plan for the day including EVA prep, research hardware set up and how best to accomplish our daily tasks. We also spent some time checking in with each crewmember to discuss factors affecting our health, stress and personal mission goals. We had a positive discussion and everyone felt heard, which is important to the leadership of this team.

We began EVA prep at 09:30 and sent 3 members of our team through the airlock to explore the Martian terrain while the 3 remaining crew set to work setting up research hardware in the greenhouse and science lab. The Artist in Residence and Crew Astronomer have been working all day to prepare a special 360 degree video of the Super Blue Blood Moon Eclipse which begins tonight. Our best viewing opportunity will begin at 04:44 am.

One very exciting event occurred after our EVA crew returned. We received a cargo shipment of our embroidered mission patches! This provided our crew with an unexpected morale boost and helped us feel that much more excited about carrying out a high fidelity analog mission worthy of an official crew mission patch.

Ad astra!
Renee Garifi
MDRS Crew 188 Executive Officer

Look Ahead Plan:
In the late evening/early morning tomorrow, all crewmembers will rise at 04:00 am to view the special astronomical mega event of the lunar eclipse of the January Blue Moon.

Anomalies in work:
MDRS Director is coordinating power system updates with contractors. Robotic observatory currently not functional. MDRS Astronomy lead working problem.

Weather:
A warm day on “Mars” with crewmembers getting quite a lot of sun on today’s EVA.

Crew Physical Status:
Crew are fighting fatigue and dehydration symptoms but are in good spirits.

EVA:
One EVA successfully completed today. We will not have an EVA tomorrow (31 January).

Reports to file:
1. Sol Summary
2. Operations Report
3. Journalist Report
4. EVA #3 Report
6. Daily Photos
7. Mission Summary
8. Green Hab Report
9. Creative Report

Support Requested: See Operations Report